OpenStudy (anonymous):

differentiate 2t^(-3/4)

OpenStudy (australopithecus):

Simple rule, f(x) = x^n f'(x) = nx^(n-1)

OpenStudy (australopithecus):

where n is just a number

OpenStudy (australopithecus):

can you solve this now?

OpenStudy (anonymous):

i will try! thanks!

OpenStudy (anonymous):

so just subtract -3/4 with one and bring that down and multiply by 2 because its 2(-3/4)?

OpenStudy (australopithecus):

yes

OpenStudy (anonymous):

thats it? so answer is -3/2t^(-7/4)?

OpenStudy (australopithecus):

This document has all the important derivative rule you need to practice and memorize, it also covers a general calc 1 courses material till the final exam.

OpenStudy (anonymous):

thank you for that document it will help since my professor just do one problem and then gives rest for homework

OpenStudy (australopithecus):

no that is wrong, f'(x) = (-3/4)2^(-3/4 - 1) f'(x) = (-3/4)2^(-3/4 - 4/4) f'(x) = (-3/4)2^(-7/4)

OpenStudy (australopithecus):

so, \[\frac{-3(2)^{-7/4}}{4}\]

OpenStudy (australopithecus):

you need to practice these rules like crazy until you can look at a function and know its derivative almost automatically

OpenStudy (anonymous):

where did the t go?

OpenStudy (anonymous):

i am lost

OpenStudy (australopithecus):

lol sorry made a mistake out of practice

OpenStudy (australopithecus):

\[\frac{(-3)(2)t^{-7/4}}{4}\]

OpenStudy (australopithecus):

the rule is very simple, just go through your homework, I cant stress practice problems enough when it comes to calculus

OpenStudy (anonymous):

i thought (-3/4)(2)t^(-7/4) would be the answer

OpenStudy (australopithecus):

yeah that is the answer

OpenStudy (australopithecus):

remember basic multiplication of fractions

OpenStudy (anonymous):

so just multiply -3/4 * 2

OpenStudy (anonymous):

(-3/2)t^(-7/4) that is correct multiplication?

OpenStudy (australopithecus):

\[\frac{-3}{4}(\frac{2t^{-7/4}}{1}) = \frac{-6t^{\frac{-7}{4}}}{4*1} = \frac{-6}{4t^{\frac{7}{4}}}\]

OpenStudy (australopithecus):

yeah that is fine i'm just crazy tired

OpenStudy (anonymous):

yeah i was like if you just multiply and stuff we get same answers you just put it in different way

OpenStudy (australopithecus):

that can be simplifed to -3/2t^(7/4)

OpenStudy (australopithecus):

sorry man losing my mind lol

OpenStudy (anonymous):

lol no problem! thanks for the help though i thought i was doing it write but i just wanted to make sure haha